# Decoding Matters: Addressing Amplification Bias and Homogeneity Issue for LLM-based Recommendation

### v1

@@ Abstract
Keqin ${}^{\spadesuit}{\,{}^{\diamondsuit}}$thanks:   Work is done during internship at Huawei Inc.     Jizhi ${}^{\spadesuit}{\,}$     Yang Zhang Xinyue ${}^{\spadesuit}{\,}$     Chong Chen    Fuli Feng University of Science and Technology of China, Huawei Inc.   National University of Singapore {baokq, cdzhangjizhi, xinyueh}@mail.ustc.edu.cn, chenchong55@huawei.com, {zyang1580, fulifeng93}@gmail.com
@@ 2.1 Large Language Models for Recommender Systems
The remarkable capabilities demonstrated by LLMs have sparked interest within the recommender systems community, leading to a surge in research aimed at integrating LLMs into these systems Wu et al. (2023); Lin et al. (2023a). This research can be broadly divided into two streams. The first stream involves directly utilizing LLMs as encoders to represent items, which are then input into conventional recommendation models Yuan et al. (2023); Xi et al. (2023). However, the prevalent LLMs are primarily decoder-only architectures, optimized for next-token-prediction, which is not conducive to information encoding. This limitation potentially constrains the full potential of LLMs in recommendation tasks.
The second stream of research focuses on harnessing the generative power of LLMs to produce recommendations directly Lin et al. (2023b); Liao et al. (2024); Wang et al. (2023); Zhang et al. (2023b, 2024). Despite this approach, it has been observed that LLMs have limited exposure to recommendation-specific data during pre-training, necessitating fine-tuning for effective application in this domain. While methods of this nature are on the rise, they often prioritize improving LLMs’ recommendation performance through training enhancements, overlooking a crucial aspect: the detailed examination of the models’ recommendation outputs and the considerations for the decoding phase.
@@ 7 Analysis
In this section, we further explore our decoding approach in detail and discuss potential extensions. First, we conduct an ablation analysis to validate the influence of our two key strategies on recommendation accuracy. We then examine the impact of the method’s designs on recommendation diversity. Finally, we investigate an extension of our methods which can enable easy adjustment of the recommendation type distribution.
@@ 8 Conclusion
In this paper, we introduce $D^{3}$ to mitigate amplification bias and the problem of homogeneity in existing LLM-based recommendation decoding strategies, without introducing new ethical dilemmas. We employ publicly available data while carefully avoiding sensitive information. Nevertheless, as recommendations depend on user behavior data, privacy issues may arise. These concerns can be managed by securing user consent. Moreover, the usage of LLMs might perpetuate unseen societal biases. We recommend comprehensive risk evaluations and cautioning users about the potential risks associated with model deployment.
@@ Appendix C Analysis of Homogeneity Issue
Figure 4 illustrates the homogeneity issue observed in the Sports, Toys, and Games datasets, a phenomenon that is consistent with findings in the other three datasets. LLM-based recommender systems demonstrate a significant level of similarity both in terms of text and category. This confirms that across all six datasets, while the integration of textual information enhances performance, it simultaneously introduces a text-level bias. This bias diminishes the diversity of the recommendations, causing them to converge on specific characteristics. It’s important to clarify that the apparent issue identified within our study is not inherently detrimental. On one side, we note that even text-free models also recommend items with textual similarities, pointing to a fundamental tendency within recommender systems. On the other side, as evidenced by Figure 6, where we have mapped out the proportion of repetitive item categories within the training set, it’s clear that the practice of suggesting items within similar categories inherently fulfills the preferences of certain user demographics. However, when deploying LLM-based recommender systems, their inherent copy mechanisms tend to exaggerate this feature, which might detract from the experience of a broader user base. Consequently, our findings advocate for a strategic adjustment to this issue, aiming for moderation rather than elimination, to optimally balance user satisfaction across the spectrum.

### v2

@@ Abstract
Keqin Bao1,2thanks:   Work is done during internship at Huawei Inc.     Jizhi Zhang1     Yang Zhang3thanks:   Corresponding Author Xinyue Huo1     Chong Chen3    Fuli Feng1† 1University of Science and Technology of China, 2Huawei Inc.   3National University of Singapore {baokq, cdzhangjizhi, xinyueh}@mail.ustc.edu.cn, chenchong55@huawei.com, {zyang1580, fulifeng93}@gmail.com
@@ 2.1 Large Language Models for Recommender Systems
The remarkable capabilities demonstrated by LLMs have sparked interest within the recommender systems community, leading to a surge in research aimed at integrating LLMs into these systems Wu et al. (2023); Lin et al. (2023a); Bao et al. (2024). This research can be broadly divided into two streams. The first stream involves directly utilizing LLMs as encoders to represent items, which are then input into conventional recommendation models Yuan et al. (2023); Xi et al. (2023). However, the prevalent LLMs are primarily decoder-only architectures, optimized for next-token-prediction, which is not conducive to information encoding. This limitation potentially constrains the full potential of LLMs in recommendation tasks.
The second stream of research focuses on harnessing the generative power of LLMs to produce recommendations directly Lin et al. (2023b); Liao et al. (2024); Wang et al. (2023); Zhang et al. (2023b, 2024a, 2024b). Despite this approach, it has been observed that LLMs have limited exposure to recommendation-specific data during pre-training, necessitating fine-tuning for effective application in this domain. While methods of this nature are on the rise, they often prioritize improving LLMs’ recommendation performance through training enhancements, overlooking a crucial aspect: the detailed examination of the models’ recommendation outputs and the considerations for the decoding phase.
2.2 Decoding in Language Model
The second stream of research focuses on harnessing the generative power of LLMs to produce recommendations directly Lin et al. (2023b); Liao et al. (2024); Wang et al. (2023); Zhang et al. (2023b, 2024b). Despite this approach, it has been observed that LLMs have limited exposure to recommendation-specific data during pre-training, necessitating fine-tuning for effective application in this domain. While methods of this nature are on the rise, they often prioritize improving LLMs’ recommendation performance through training enhancements, overlooking a crucial aspect: the detailed examination of the models’ recommendation outputs and the considerations for the decoding phase.
@@ 7 Analysis
This section further explores our decoding approach in detail and discusses potential extensions. First, we conduct an ablation analysis to validate the influence of our two key strategies on recommendation accuracy. We then examine the impact of the method’s designs on recommendation diversity and investigate an extension of our methods which can enable easy adjustment of the recommendation type distribution. Finally, we validate the generalizability of the proposed method using additional backbones, approaches, and datasets.
@@ 7.3 Distribution Adjustment
Datasets CID +$D^{3}$ BIGRec +$D^{3}$
CDs 0.0610 0.0921 0.0869 0.1179
Games 0.0447 0.0598 0.0727 0.0856
Table 3: We report the HR@10 results of our methods on another approach – CID Hua et al. (2023) and another backbone –Llama3.1-8B Dubey et al. (2024). “+$D^{3}$” denotes applying our decoding method. The best results are bolded.
@@
7.4 Generalizability
To show the generalizability of our methods, we conduct further experiments using additional backbones, approaches, and datasets. In detail, as shown in Table 3, we present the results of applying our decoding method to a new backbone model, Llama3.1-8B Dubey et al. (2024), and a new item representation approach, CID Hua et al. (2023), on the CDs and Games datasets.
Metrics BIGRec +$D^{3}$
NDCG@10 0.0277 0.0398
HR@10 0.0563 0.0814
Table 4: We report the NDCG@10 and HR@10 results of our methods on steam datasets Kang and McAuley (2018). “+$D^{3}$” denotes applying our decoding method. The best results are bolded.
The findings demonstrate consistent improvements in performance with both the new backbone and the new method, highlighting the generalizability and effectiveness of our decoding approach. Besides, we have also applied our approach to the Steam Kang and McAuley (2018) dataset. Given the substantial size of this dataset, we randomly selected 10% of the users to expedite the validation process of our method. The results of our experiments are presented in Table  4. These findings demonstrate that our decoding method is not only applicable to other platforms and datasets but also further validates the generalizability of our approach.
@@ 8 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62272437), and the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Acknowledgments
In this paper, we introduce $D^{3}$ to mitigate amplification bias and the problem of homogeneity in existing LLM-based recommendation decoding strategies, without introducing new ethical dilemmas. We employ publicly available data while carefully avoiding sensitive information. Nevertheless, as recommendations depend on user behavior data, privacy issues may arise. These concerns can be managed by securing user consent. Moreover, the usage of LLMs might perpetuate unseen societal biases. We recommend comprehensive risk evaluations and caution users about the potential risks associated with model deployment.
@@ Appendix C Analysis of Homogeneity Issue
Figure 4 illustrates the homogeneity issue observed in the Sports, Toys, and Games datasets, a phenomenon that is consistent with findings in the other three datasets. LLM-based recommender systems demonstrate a significant level of similarity both in terms of text and category. This confirms that across all six datasets, while the integration of textual information enhances performance, it simultaneously introduces a text-level bias. This bias diminishes the diversity of the recommendations, causing them to converge on specific characteristics. It’s important to clarify that the apparent issue identified within our study is not inherently detrimental. On one side, we note that even text-free models also recommend items with textual similarities, pointing to a fundamental tendency within recommender systems. On the other side, as evidenced by Figure 6, where we have mapped out the proportion of repetitive item categories within the training set, it’s clear that the practice of suggesting items within similar categories inherently fulfills the preferences of certain user demographics. However, when deploying LLM-based recommender systems, their inherent copy mechanisms tend to exaggerate this feature, which might detract from the experience of a broader user base. Consequently, our findings advocate for a strategic adjustment to this issue, aiming for moderation rather than elimination, to balance user satisfaction across the spectrum optimally.